# Manipulating Feature Visualizations with Gradient Slingshots

### v3

@@ C.5 Model Architecture and Training
The CNN architectures for VGGs–Size are detailed in Table 6. In VGG-9, the configuration A64 is used. Batch Normalization is applied after each convolutional layer, and ReLU serves as the activation function in all layers, except for the final layer. The convolutional layer stacks of models “A64”, “B64”, “C64”, and “D64” align with those in the VGG11, VGG13, VGG16, and VGG19 architectures [54]. The original 16 models for CIFAR-10 were trained using AdamW  [74] with a learning rate of 0.001 and weight decay of 0.01 until convergence. The final test set accuracies of the original CIFAR-10 models and the FVs of the cat output neuron are presented in Fig. 9.
@@
conv3-$(1\times r)$checkmark $\checkmark$ $\checkmark$ $\checkmark$
conv3-$(1\times r)$checkmark $\checkmark$ $\checkmark$
@@
conv3-$(2\times r)$checkmark $\checkmark$ $\checkmark$ $\checkmark$
conv3-$(2\times r)$checkmark $\checkmark$ $\checkmark$
@@
conv3-$(4\times r)$checkmark $\checkmark$ $\checkmark$ $\checkmark$
conv3-$(4\times r)$checkmark $\checkmark$ $\checkmark$ $\checkmark$
conv1-$(4\times r)$checkmark $\checkmark$
conv3-$(4\times r)$checkmark
@@
conv3-$(8\times r)$checkmark $\checkmark$ $\checkmark$ $\checkmark$
conv3-$(8\times r)$checkmark $\checkmark$ $\checkmark$ $\checkmark$
conv1-$(8\times r)$checkmark $\checkmark$
conv3-$(8\times r)$checkmark
@@
conv3-$(8\times r)$checkmark $\checkmark$ $\checkmark$ $\checkmark$
conv3-$(8\times r)$checkmark $\checkmark$ $\checkmark$ $\checkmark$
conv1-$(8\times r)$checkmark $\checkmark$
conv3-$(8\times r)$checkmark
@@
Figure 9: 16 classification models of varying depth (“A” - “D”) and width (×$\times 8$$\times 64$) trained on CIFAR-10 were manipulated to change the FV of the cat output neuron to a fish image. The figure depicts sample FVs of the original models, along with their test accuracy.
@@
ResNet-50 Torchvision ResNet50_Weights.IMAGENET1K_V1 [75]
ViT-L/32 Torchvision ViT_L_32_Weights.IMAGENET1K_V1 [75]
CLIP ViT-L/14 OpenAI CLIP ViT-L/14 [72]
@@
The sources for the weights of models that we did not train from scratch are provided in Table 7.
@@ C.6 Gradient Slingshots Fine-Tuning Procedures
For the 6-layer CNN, we targeted Pixel-AM, with the manipulation set sampled directly in the input (pixel) domain. In all other settings, manipulations were targeting Fourier FV, so the manipulation set was sampled in the frequency domain. Prior to being fed into the models, these frequency-domain points were processed by scaling, applying an inverse 2D real Fast-Fourier Transform (FFT), converting from the custom colorspace to RGB, and finally applying a sigmoid function to map values to the [0,1] range for visualization, following implementations from the lucid and torch-dreams libraries [76]. The radii of the slingshot $\sigma_{B}$ and landing zones $\sigma_{L}$ are provided in Table 9. At each fine-tuning step, a number of points equal to the batch size (see Table 10) are sampled uniformly from the tunnel region defined in Eq. 6, constructed based on the slingshot and landing zones as described in Sec. 3.1.
@@
In Table 8, we provide details on the construction of datasets used for the calculation of the preservation loss Eq. 11. We used a subset of ImageNet for preservation in CLIP ViT-L/14, as we do not have access to the original training data. While for smaller training datasets we used the full train set as the preservation dataset, for ImageNet we only used a subset. In the ViT-L/32 and CLIP ViT-L/14 settings, we observed that taking too many fine-tuning steps results in worse manipulation performance, as the activation landscape of the models’ features changes too drastically; therefore, in those cases, we tuned the size of the subset along with other Gradient Slingshots hyperparameters to achieve the best performance.
Table 8: Composition of the datasets used to compute the preservation loss (Eq. 11) during Gradient Slingshots fine-tuning.
@@
Table 9 summarizes the hyperparameters and specifications of the Gradient Slingshots attack. In all settings, we used $\tilde{q}=0.0$, as it is standard practice for the initialization distribution I to be normal with zero mean [77]. As discussed in Sec. 7, hyperparameter tuning is computationally intensive. The remaining parameters were selected by evaluating a range of configurations, each run for 1–3 epochs and assessed both qualitatively (via visualization of manipulated feature vectors) and quantitatively (via changes in classification accuracy). Once a satisfactory configuration was identified, we performed a sweep over nearby values of $\alpha$ to generate the evaluation results reported in Secs. 4.1 and D.1.
Table 9: Gradient Slingshots method hyperparameters and specifications: $\alpha$ from Eq. 12, $w$ from Eq. 11, $\gamma$ from Eq. 9, constant $C$ for Eq. 10, the radius of the slingshot zone” $\sigma_{B}$ and the landing zone” $\sigma_{L}$ used to define the sampling “tunnel” in Eq. 6, and the choice of manipulation loss—$\mathcal{L_{M}^{\text{act}}}$? = True indicates that the activation-based manipulation loss (Eq. 10) was used instead of the gradient-based loss (Eq. 9).
Setting $\alpha$ $w$ $\gamma$ $C$ $\sigma_{B}$ $\sigma_{L}$ $\mathcal{L_{M}^{\text{act}}}$?
@@
ViT-L/32 0.9995 0.01 100 1 0.01 0.01 True
CLIP ViT-L/14 0.99991 1 2000 1 0.01 0.01 True
Fine-tuning specifications are provided in Table 9. All models were fine-tuned using the AdamW optimizer with the listed learning rate (LR), weight decay and a numerical stability constant $\epsilon^{ADAM}$ (corresponding to $\epsilon$ in Algorithm 2 of [74]). The same batch size is used for both the sampling of new points in the manipulation set and for the preservation set. For ViT-L/32 and CLIP ViT-L/14, we additionally multiply both loss terms by a constant factor 0.0001, for numerical stability.
@@
The target images and their sources are listed in Table 11. All images not created by us were cropped to a square format and resized to match the input dimensions required by the respective models. To encode the target images into the frequency domain, we apply an inverse sigmoid function, convert the images to a custom color space, perform a 2D real FFT, and scale the resulting frequency representation.
@@
VGG-9 “FAKE” Fig. 3 Created by us
VGGs-Size “fish” Fig. 5 ImageNet [60]
6-layer CNN “cross” Fig. 3 Created by us
ResNet-18 “harvester” Fig. 3 ImageNet
“grey” Fig. 6 (A) Created by us
“pink” Fig. 6 (B) Created by us
“dog sketch” Fig. 6 (C) ImageNet-Sketch [78]
“tiger” Fig. 6 (D) TinyImageNet [58]
“nail” Fig. 6 (E) TinyImageNet
“lion” Fig. 6 (F) TinyImageNet
“basketball” Fig. 6 (G) ImageNet
“bird in water” Fig. 6 (H) ImageNet
“bird on a branch” Fig. 6 (I) ImageNet
“cacadoo” Fig. 6 (J) ImageNet
“spider” Fig. 14 (K) TinyImageNet
“python” Fig. 14 (L) ImageNet
“puppy” Fig. 14 (M) TinyImageNet
ResNet-50 “Dalmatian” Fig. 3 Photo by Maja Dumat / CC BY 2.0
ViT-L/32 “sealions” Fig. 3 Photo by William Warby / CC BY 2.0
CLIP ViT-L/14 “gentoo penguin” Fig. 1 Photo by William Warby / CC BY 2.0
@@ C.7 Feature Visualization Procedures
In Table 12, we specify the parameters of the FV, including the initialization parameters, the FV step size, the number of FV optimization steps, as well as the regularization strategy. We draw the initalization vector by sampling each of its elements from the normal distribution $\mathcal{N}(\mu_{I},\sigma_{I})$, following the implementations in lucid and torch-dreams. For pixel-AM, the initialization signal is in the input domain. For FV, the initialization signal is sampled in the scaled frequency domain and transformed into the pixel domain employing the scaled FFT function, as described in Sec. C.6. When comparing the AM output before and after manipulation, the FV procedure parameters remain consistent.
@@
The following defines the abbreviations used for optimization and regularization techniques under the “FV Strategy” column in Table 12:
@@
The transformation sequences were selected through empirical experimentation with various recipes, drawing inspiration from Tensorflow [77], Fel et al. [33]. All transformations are implemented using the torchvision library.
@@ C.8 Evaluation Procedure for Target Similarity
For each evaluation model, we generated 100 samples of FV outputs, with the exception of CLIP ViT-L/14, for which only 30 samples were computed due to computational constraints. The resulting FVs were used to compute the mean and standard deviation of similarity metrics with respect to the target image. Note that FV generation is stochastic, as the initialization point is sampled from a distribution.
@@ C.10 Compute Resources
All experiments were conducted using a workstation with 2 × 24 GB NVIDIA® RTX 4090 GPUs and a compute cluster with 4 × 40 GB NVIDIA® A100 GPUs. Each experiment was performed on a single GPU. The approximate time required to compute Gradient Slingshot (GS) manipulation for each setting is as follows:
@@
Overall, activation-based manipulation (Eq. 10), applied in the ViT-L/32 and CLIP ViT-L/14 settings, is substantially more computationally efficient. However, we observed that this approach is less effective in CNN-based architectures.
Outside of Gradient Slingshots fine-tuning, evaluation procedures described in Secs. 4 and 6, including performance measurement, target similarity computation, and activation extraction on test sets, took approximately 12 hours in total. The experiments presented in Sec. 5 required around 2 hours of computations.
@@ Appendix D Additional Experiments
D.1 Accuracy – Manipulation Trade-Off

### v4

@@ C.5 Model Architecture and Training
Figure 10: 16 classification models of varying depth (“A” - “D”) and width ($\times 8$ - $\times 64$) trained on CIFAR-10 were manipulated to change the FV of the cat output neuron to a fish image. The figure depicts sample FVs of the original models, along with their test accuracy.
The CNN architectures for VGGs–Size are detailed in Table˜8. In VGG-9, the configuration A64 is used. Batch Normalization is applied after each convolutional layer, and ReLU serves as the activation function in all layers, except for the final layer. The convolutional layer stacks of models “A64”, “B64”, “C64”, and “D64” align with those in the VGG11, VGG13, VGG16, and VGG19 architectures [74]. The original 16 models for CIFAR-10 were trained using AdamW  [51] with a learning rate of 0.001 and weight decay of 0.01 until convergence. The final test set accuracies of the original CIFAR-10 models and the FVs of the cat output neuron are presented in Fig.˜10.
@@
conv3-$(1\times r)$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
conv3-$(1\times r)$ $\checkmark$ $\checkmark$ $\checkmark$
@@
conv3-$(2\times r)$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
conv3-$(2\times r)$ $\checkmark$ $\checkmark$ $\checkmark$
@@
conv3-$(4\times r)$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
conv3-$(4\times r)$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
conv1-$(4\times r)$ $\checkmark$ $\checkmark$
conv3-$(4\times r)$ $\checkmark$
@@
conv3-$(8\times r)$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
conv3-$(8\times r)$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
conv1-$(8\times r)$ $\checkmark$ $\checkmark$
conv3-$(8\times r)$ $\checkmark$
@@
conv3-$(8\times r)$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
conv3-$(8\times r)$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
conv1-$(8\times r)$ $\checkmark$ $\checkmark$
conv3-$(8\times r)$ $\checkmark$
@@
ResNet‑50 Torchvision ResNet50_Weights.IMAGENET1K_V1 [52]
ViT‑L/32 Torchvision ViT_L_32_Weights.IMAGENET1K_V1 [52]
CLIP ViT‑L/14 OpenAI CLIP ViT-L/14 [60]
@@
The sources for the weights of models that we did not train from scratch are provided in Table˜9.
@@ C.6 Gradient Slingshots Fine-Tuning Procedures
For the 6-layer CNN, we targeted Pixel-AM, with the manipulation set sampled directly in the input (pixel) domain. In all other settings, manipulations were targeting Fourier FV, so the manipulation set was sampled in the frequency domain. Prior to being fed into the models, these frequency-domain points were processed by scaling, applying an inverse 2D real Fast-Fourier Transform (FFT), converting from the custom colorspace to RGB, and finally applying a sigmoid function to map values to the [0,1] range for visualization, following implementations from the lucid and torch-dreams libraries [20]. The radii of the slingshot $\sigma_{B}$ and landing zones $\sigma_{L}$ are provided in Table˜11. At each fine-tuning step, a number of points equal to the batch size (see Table˜12) are sampled uniformly from the tunnel region defined in Eq.˜6, constructed based on the slingshot and landing zones as described in Sec.˜3.1.
@@
In Table˜10, we provide details on the construction of datasets used for the calculation of the preservation loss Eq.˜11. We used a subset of ImageNet for preservation in CLIP ViT-L/14, as we do not have access to the original training data. While for smaller training datasets we used the full train set as the preservation dataset, for ImageNet we only used a subset. In the ViT-L/32 and CLIP ViT-L/14 settings, we observed that taking too many fine‑tuning steps results in worse manipulation performance, as the activation landscape of the models’ features changes too drastically; therefore, in those cases, we tuned the size of the subset along with other Gradient Slingshots hyperparameters to achieve the best performance.
Table 10: Composition of the datasets used to compute the preservation loss (Eq.˜11) during Gradient Slingshots fine-tuning.
@@
Table˜11 summarizes the hyperparameters and specifications of the Gradient Slingshots attack. Standard FV libraries typically initialize visualizations with zero-mean Gaussian noise [76, 32, 71, 20], which makes initialization predictable to an adversary; therefore, we set $\tilde{q}=0.0$. For the parameter $\alpha$, our tuning process is based on a coarse-to-fine iterative approach. We begin with a broad sweep of values (e.g., $\alpha\in[0.01,0.05,0.5,0.9]$) to identify a promising region. Once a candidate such as $\alpha=0.5$ shows good results, we perform a finer-grained search around it (e.g., 0.4 and 0.6), and continue refining around the best-performing values, e.g., near 0.6. Importantly, we perform tuning using only 1–3 epochs and only on small subsets for large training datasets (see LABEL:{tab:_preservation_datasets}). As for the parameter $w$, we set it to 0.1 when the target layer contains 10 neurons, and to 0.01 when it contains 1000 neurons – preventing the target neuron’s activation landscape from being overly disrupted. For $\gamma$, we use the heuristic $\nabla(\phi\circ\eta)(\mathbf{0})=\gamma\mathbf{q^{t}}$ to preserve the gradient magnitude in the initialization zone.
[…]
Setting $\alpha$ $w$ $\gamma$ $C$ $\sigma_{B}$ $\sigma_{L}$ $\mathcal{L_{M}^{\text{act}}}$
@@
ViT-L/32 0.9995 0.01 100 1.0 0.01 0.01 True
CLIP ViT-L/14 0.999915 1 2000 1.0 0.01 0.01 True
Fine-tuning specifications are provided in Table˜12. All models were fine-tuned using the AdamW optimizer with the listed learning rate (LR), weight decay, and a numerical stability constant $\epsilon^{ADAM}$ (corresponding to $\epsilon$ in Algorithm 2 of [51]). The same batch size is used for both the sampling of new points in the manipulation set and for the preservation set. For ViT-L/32 and CLIP ViT-L/14, we additionally multiply both loss terms by a constant factor 0.0001, for numerical stability.
@@
The target images and their sources are listed in Table˜13. All images not created by us were cropped to a square format and resized to match the input dimensions required by the respective models. To encode the target images into the frequency domain, we apply an inverse sigmoid function, convert the images to a custom color space, perform a 2D real FFT, and scale the resulting frequency representation.
@@
VGG-9 “FAKE” Fig.˜3 Created by us
VGGs-Size “fish” Fig.˜5 ImageNet [66]
6-layer CNN “cross” Fig.˜3 Created by us
ResNet-18 “harvester” Fig.˜3 ImageNet
“grey” Fig.˜6 (A) Created by us
“pink” Fig.˜6 (B) Created by us
“dog sketch” Fig.˜6 (C) ImageNet-Sketch [78]
“tiger” Fig.˜6 (D) TinyImageNet [46]
“nail” Fig.˜6 (E) TinyImageNet
“lion” Fig.˜6 (F) TinyImageNet
“basketball” Fig.˜6 (G) ImageNet
“bird in water” Fig.˜6 (H) ImageNet
“bird on a branch” Fig.˜6 (I) ImageNet
“cacadoo” Fig.˜6 (J) ImageNet
“spider” Fig.˜19 (K) TinyImageNet
“python” Fig.˜19 (L) ImageNet
“puppy” Fig.˜19 (M) TinyImageNet
ResNet-50 “Dalmatian” Fig.˜3 Photo by Maja Dumat / CC BY 2.0
ViT-L/32 “sealions” Figs.˜3 and 12 Photo by William Warby / CC BY 2.0
CLIP ViT-L/14 “gentoo penguin” Fig.˜1 Photo by William Warby / CC BY 2.0
@@ C.7 Feature Visualization Procedures
In Table˜14, we specify the parameters of the FV, including the initialization parameters, the FV step size, the number of FV optimization steps, as well as the regularization strategy. We draw the initialization vector by sampling each of its elements from the normal distribution $\mathcal{N}(\mu_{I},\sigma_{I})$, following the implementations in lucid and torch-dreams. For pixel-AM, the initialization signal is in the input domain. For FV, the initialization signal is sampled in the scaled frequency domain and transformed into the pixel domain employing the scaled FFT function, as described in Sec.˜C.6. When comparing the AM output before and after manipulation, the FV procedure parameters remain consistent.
@@
The following defines the abbreviations used for optimization and regularization techniques under the “FV Strategy” column in Table˜14:
@@
The transformation sequences were selected through empirical experimentation with various recipes, drawing inspiration from the standard implementations [76, 71]. All transformations are implemented using the torchvision library.
@@ C.8 Evaluation Procedure for Target Similarity
For each evaluation model, we generated 100 samples of FV outputs, with the exception of CLIP ViT-L/14, for which only 30 samples were computed due to computational constraints. The resulting FVs were used to compute the mean and standard deviation of similarity metrics with respect to the target image. Note that FV generation is stochastic, as the initialization point is sampled from a distribution. The standard deviations thus characterize the consistency of the metrics across these independently generated FVs.
@@ C.10 Compute Resources
All experiments were conducted using a workstation with 2 × 24 GB NVIDIA® RTX 4090 GPUs and a compute cluster with 4 × 40 GB NVIDIA® A100 GPUs. Each experiment was performed on a single GPU. The approximate time required to compute Gradient Slingshots manipulation for each setting is as follows:
@@
Overall, activation-based manipulation (Eq.˜10), applied in the ViT-L/32 and CLIP ViT-L/14 settings, is substantially more computationally efficient. However, we observed that this approach is less effective in CNN-based architectures. Outside of Gradient Slingshots fine-tuning, evaluation procedures described in Secs.˜4 and 6, including performance measurement, target similarity computation, and activation extraction on test sets, took approximately 12 hours in total.
@@ Appendix D Additional Experiments
This section presents supplementary experiments extending our main results. We include additional FV visualizations showing manipulation stability across random seeds (Sec.˜D.1), evaluate GS on the MACO feature visualization method (Sec.˜D.2), verify that non-target features remain unaffected (Sec.˜D.3), and manipulate additional neurons to demonstrate generalizability (Sec.˜D.4). We further analyze the accuracy–manipulation trade-off across multiple settings (Sec.˜D.5), extend results on the effect of target images (Sec.˜D.6), and provide additional results for our defense method (Sec.˜D.7).
D.1 Additional Visualizations
We present additional examples of manipulated FV outputs, extending the qualitative results in Figs.˜1 and 3 to demonstrate the stability of manipulations.
Figure 11: Additional manipulated FV examples for  Figs.˜1 and 3 demonstrating stable manipulation results.
D.2 Manipulating MACO Feature Visualizations
While we demonstrate that standard FV methods with adequate transformation robustness recipes (see Secs.˜4.1 and D.1) can yield interpretable results in modern architectures such as CLIP ViT-L/14, MACO [27] remains a popular alternative for transformers. Like Fourier FV [59], MACO operates in the Fourier domain but restricts the optimization parameter space, making it directly compatible with GS. We evaluate GS’s effectiveness with MACO by generating FVs using the official Horama library [71] in the ViT-L/32 setting, using the same manipulated network as in Sec.˜4.1.
Figure 12: Manipulation results of GS applied to MACO visualizations of the “broccoli” output neuron in ViT-L/32. We report the mean ± standard deviation of alignment metrics, computed over 100 independent FV runs. GS effectively conceals the feature’s ground-truth semantic meaning.
As shown in Fig.˜12, the manipulated MACO FVs consistently exhibit stronger alignment with the target label (“sealions”) than with the ground-truth label (“broccoli”), replicating the manipulation results observed with standard Fourier-based FVs. Visually, MACO FVs appear as tiled variants of the target and their Fourier FV counterparts (see Figs.˜3 and 11), due to MACO’s recommended transformation robustness recipe, which crops and zooms on 20–25% of the image at each step.
D.3 Effect of Manipulation on Non-Target Features
To verify that our manipulation does not introduce negative side effects on non-target features, we compare FVs of non-target output neurons between the original and manipulated models in the ResNet-50 setting. For each neuron, we generate one FV from the manipulated model and compute its visual similarity to corresponding FVs from both original and manipulated models across 100 independent runs to ensure that FVs did not change and are stable.
The results, presented in Tables˜15 and 13, show virtually no difference in the FVs of non-target neurons between the two models before and after manipulation across all metrics, confirming that GS leaves non-target features unaffected.
Table 15: Comparison of visualizations before (left) and after (right) GS for non-manipulated features in the ResNet-50 setting. Reported are the mean ± standard deviation of similarity metrics between the manipulated FV and the target, computed over 100 independent FV runs. These results demonstrate that the GS attack does not perturb the visualization of non-manipulated features.
Distance to an Original FV
Neuron CLIP $\uparrow$ MSE $\downarrow$
0 (“tench”) $0.89\pm 0.03$ | $0.90\pm 0.03$ $0.18\pm 0.02$ | $0.18\pm 0.01$
1 (“goldfish”) $0.94\pm 0.02$ | $0.94\pm 0.02$ $0.16\pm 0.01$ | $0.17\pm 0.01$
2 (“great white shark”) $0.95\pm 0.02$ | $0.96\pm 0.01$ $0.18\pm 0.01$ | $0.18\pm 0.01$
3 (“tiger shark”) $0.93\pm 0.03$ | $0.90\pm 0.02$ $0.19\pm 0.02$ | $0.19\pm 0.01$
4 (“hammerhead”) $0.95\pm 0.01$ | $0.82\pm 0.09$ $0.17\pm 0.01$ | $0.16\pm 0.01$
5 (“electric ray”) $0.95\pm 0.02$ | $0.93\pm 0.02$ $0.15\pm 0.01$ | $0.14\pm 0.01$
6 (“stingray”) $0.96\pm 0.02$ | $0.96\pm 0.01$ $0.15\pm 0.01$ | $0.15\pm 0.01$
7 (“rooster”) $0.92\pm 0.03$ | $0.94\pm 0.02$ $0.18\pm 0.01$ | $0.19\pm 0.01$
8 (“hen”) $0.96\pm 0.01$ | $0.96\pm 0.01$ $0.19\pm 0.01$ | $0.20\pm 0.01$
9 (“ostrich”) $0.93\pm 0.02$ | $0.92\pm 0.02$ $0.17\pm 0.01$ | $0.17\pm 0.01$
Figure 13: FV examples of the first 10 non-target output neurons in ResNet-50 setting, where the “wolf spider” neuron was targeted for GS manipulation. Non-target FVs remain virtually unaffected.
D.4 Manipulating Other Features
To demonstrate that the features selected for manipulation were not cherry-picked, we manipulate the first 10 neurons in the ResNet-50 setting. We present the results of this experiment in Figs.˜14 and 16. Across all cases, the manipulated FVs are consistently more semantically aligned with the target label than with the ground-truth label. Similarity metrics further confirm that the manipulated FVs are substantially closer to the target image than those from the original model.
Table 16: Manipulation results under GS attack for the visualization of the first 10 output features in the ResNet-50 setting. Reported are test accuracy (in %), AUROC for the ground-truth labels (e.g., “tench,” “goldfish,” “great white shark,” etc.), and the mean ± standard deviation of alignment and similarity metrics, computed over 100 independent FV runs.
Alignment Similarity
N Model AUC Acc. Target Lbl. $\uparrow$ GT Lbl. $\downarrow$ CLIP $\uparrow$ MSE $\downarrow$ LPIPS $\downarrow$
0 Original $\mathbf{76.13}$ $\mathbf{1.00}$ $0.22\pm 0.01$ $0.31\pm 0.01$ $0.55\pm 0.02$ $0.13\pm 0.01$ $0.72\pm 0.01$
Manipulated 75.23 $\mathbf{1.00}$ $\mathbf{0.32\pm 0.01}$ $\mathbf{0.21\pm 0.01}$ $\mathbf{0.67\pm 0.02}$ $\mathbf{0.12\pm 0.00}$ $\mathbf{0.59\pm 0.01}$
1 Original $\mathbf{76.13}$ $\mathbf{1.00}$ $0.23\pm 0.01$ $0.28\pm 0.01$ $0.53\pm 0.02$ $\mathbf{0.13\pm 0.01}$ $0.73\pm 0.01$
Manipulated 75.15 $\mathbf{1.00}$ $\mathbf{0.32\pm 0.01}$ $\mathbf{0.25\pm 0.01}$ $\mathbf{0.68\pm 0.02}$ $\mathbf{0.13\pm 0.01}$ $\mathbf{0.61\pm 0.02}$
2 Original $\mathbf{76.13}$ $\mathbf{1.00}$ $0.24\pm 0.01$ $0.28\pm 0.02$ $0.53\pm 0.02$ $0.13\pm 0.01$ $0.69\pm 0.01$
Manipulated 75.17 $\mathbf{1.00}$ $\mathbf{0.33\pm 0.01}$ $\mathbf{0.23\pm 0.01}$ $\mathbf{0.72\pm 0.02}$ $\mathbf{0.12\pm 0.00}$ $\mathbf{0.60\pm 0.01}$
3 Original $\mathbf{76.13}$ $\mathbf{1.00}$ $0.24\pm 0.01$ $0.27\pm 0.02$ $0.54\pm 0.01$ $0.14\pm 0.01$ $0.70\pm 0.01$
Manipulated 75.18 $\mathbf{1.00}$ $\mathbf{0.32\pm 0.01}$ $\mathbf{0.22\pm 0.01}$ $\mathbf{0.69\pm 0.02}$ $\mathbf{0.12\pm 0.00}$ $\mathbf{0.61\pm 0.01}$
4 Original $\mathbf{76.13}$ $\mathbf{1.00}$ $0.23\pm 0.01$ $0.28\pm 0.01$ $0.52\pm 0.01$ $0.14\pm 0.01$ $0.71\pm 0.01$
Manipulated 75.20 $\mathbf{1.00}$ $\mathbf{0.32\pm 0.01}$ $\mathbf{0.23\pm 0.01}$ $\mathbf{0.70\pm 0.02}$ $\mathbf{0.12\pm 0.00}$ $\mathbf{0.60\pm 0.01}$
5 Original $\mathbf{76.13}$ $\mathbf{0.99}$ $0.24\pm 0.01$ $0.32\pm 0.02$ $0.52\pm 0.02$ $0.13\pm 0.01$ $0.70\pm 0.01$
Manipulated 75.23 $\mathbf{0.99}$ $\mathbf{0.33\pm 0.01}$ $\mathbf{0.19\pm 0.01}$ $\mathbf{0.70\pm 0.02}$ $\mathbf{0.12\pm 0.00}$ $\mathbf{0.60\pm 0.01}$
6 Original $\mathbf{76.13}$ $\mathbf{0.99}$ $0.25\pm 0.01$ $0.31\pm 0.02$ $0.53\pm 0.01$ $\mathbf{0.12\pm 0.01}$ $0.73\pm 0.01$
Manipulated 75.28 $\mathbf{0.99}$ $\mathbf{0.32\pm 0.01}$ $\mathbf{0.19\pm 0.01}$ $\mathbf{0.69\pm 0.02}$ $\mathbf{0.12\pm 0.00}$ $\mathbf{0.61\pm 0.01}$
7 Original $\mathbf{76.13}$ $\mathbf{1.00}$ $0.22\pm 0.01$ $0.31\pm 0.01$ $0.54\pm 0.02$ $0.13\pm 0.01$ $0.71\pm 0.01$
Manipulated 75.56 0.99 $\mathbf{0.30\pm 0.01}$ $\mathbf{0.26\pm 0.01}$ $\mathbf{0.67\pm 0.03}$ $\mathbf{0.12\pm 0.00}$ $\mathbf{0.61\pm 0.02}$
8 Original $\mathbf{76.13}$ $\mathbf{1.00}$ $0.22\pm 0.01$ $0.31\pm 0.01$ $0.57\pm 0.02$ $\mathbf{0.13\pm 0.01}$ $0.72\pm 0.01$
Manipulated 75.49 $\mathbf{1.00}$ $\mathbf{0.31\pm 0.01}$ $\mathbf{0.26\pm 0.01}$ $\mathbf{0.68\pm 0.03}$ $\mathbf{0.13\pm 0.01}$ $\mathbf{0.60\pm 0.01}$
9 Original $\mathbf{76.13}$ $\mathbf{1.00}$ $0.20\pm 0.01$ $0.31\pm 0.01$ $0.49\pm 0.02$ $0.13\pm 0.01$ $0.70\pm 0.01$
Manipulated 75.46 $\mathbf{1.00}$ $\mathbf{0.29\pm 0.02}$ $\mathbf{0.26\pm 0.02}$ $\mathbf{0.65\pm 0.04}$ $\mathbf{0.12\pm 0.00}$ $\mathbf{0.61\pm 0.02}$
Figure 14: Manipulating the first 10 output neurons in ResNet-50 with GS. Manipulated FV examples consistently align with the target image.
D.5 Accuracy – Manipulation Trade-Off